A =

100

1. Compute all four spaces of the matrix .

2. Find bases in them.

3. Find the matrix of A|c(A^T) in the bases you computed
Answer #1

1). The four fundamental subspaces are the column space and the null space of A and AT. Also, Col (AT) = Null(A) and Null(AT) = Col(A). To determine these 4 subspaces, we will reduce A to its RREF as under:

1. Add -2 times the 2nd row to the 3rd row

2. Add -1 times the 2nd row to the 1st row

Then the RREF of A is

1

0

0

0

1

1

0

0

0

This implies that the first 2 columns of A are linearly independent and its 3rd column is same as its 2nd column. Therefore, Col(A) = span {(1,0,0)T,(1,1,2)T}.

Null(A) is the set of solutions to the equation AX = 0. If X = (x,y,z)T, then the equation AX= 0 is equivalent to x = 0 and y+z = 0 or, y = -z. Then X = (0,-z,z)T = z(0,-1,1)T. This implies that every solution to the the equation AX= 0 is a scalar multiple of the vector(0,-1,1)T. Hence, Null(A) = span {(0,-1,1)T}.

Now, let X = (x,y,z)T be an arbitrary vector in Col (AT) = Null(A).

Then (x,y,z)T. (0,-1,1)T = 0 or, -y+z = 0 or, y = z. Then X = (x,z,z)T = x(1,0,0)T+ z( 0,1,1)T. This implies that every vector in Col (AT) = Null(A) is a linear combination of 2 linearly independent vectors (1,0,0)T,( 0,1,1)T. Hence, Col (AT) = Null(A) = span {(1,0,0)T,( 0,1,1)T }.

Further, let X = (x,y,z)T be an arbitrary vector in Null(AT) = Col(A).

Then (x,y,z)T. (1,0,0)T = 0 or, x = 0 and (x,y,z)T. (1,1,2)T = 0 or,x+y+2z = 0 or, y+2z = 0 or, y = -2z. Then X = (0,-2z,z)T = z(0,-2,1)T. This implies that every vector in Null(AT) is a scalar multiple of the vector(0,-2,1)T. Hence, Null(AT)   = Col(A) =span{(0,-2,1)T }.

2). The set {(1,0,0)T,(1,1,2)T} is a basis for Col(A).

The set {(0,-1,1)T} is a basis for Null(A).

The set {(1,0,0)T,( 0,1,1)T } is a basis for Col (AT) = Null(A).

The set {(0,-2,1)T } is a basis for Null(AT) = Col(A).

3). Please advise the meaning of the notation A|c(A^T).